OpenStudy (anonymous):

Describe the transformations required to obtain the graph of the function f(x) from the graph of the function g(x). f(x)= cos x/4 g(x)= cos x

OpenStudy (campbell_st):

if you have cos(bx) then the period is \[\frac{2\pi}{b}\] in the 1st curve f(x) it takes a distance of \[\frac{2\pi}{1}\] to complete 1 full cycle of the curve, this is known as the period. in g(x) the value of b is 1/4 so the period is \[\frac{2\pi}{\frac{1}{4}} = 8\pi\] or the cycle of g(x) is 4 times longer than the cycle for g(x) to change f(x) to g(x) the value of b, in f(x) is divided by 4 hope this makes sense.

OpenStudy (jdoe0001):

\(\large {\textit{parent function } cos(x)\qquad \begin{array}{cccllll} g(x)=&cos\left(\frac{x}{\color{red}{4}}\right)\\ &\qquad \uparrow\\ \end{array}\\ \quad \\ \bf \textit{period change from regular period of }2\pi\\ \quad \\ \textit{to new period of }\cfrac{2\pi}{\color{red}{4}}}\)

OpenStudy (anonymous):

is it a horizontal or vertical shift of 1/4

OpenStudy (campbell_st):

there is no shift..... the curve g(x) is stretched by a factor of 4 have a look at the graphs g(x) takes a distance of 8pi to complete a cycle... and f(x) takes a distance of 2pi or f(x) completes 4 cycles over the same distance it takes g(x) to complete 1 cycle